# Exponentially accelerated relaxation and quantum Mpemba effect in open quantum systems

### v1

@@ V Examples
Figure 3: (Color online) The quantum Mpemba effect in the dynamics of the XXZ model with open boundary conditions [see details in Sec. V], where $N=5$ spins, $\Delta=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous state $\rho(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
Figures 2 and 3 present the numerical simulations for TFI and XXZ models, respectively. We set $N=5$, $h=J/2$, $\Delta=J/2$, $\gamma=1$, $T=0.1$, and ${k_{B}}=1$. In these figures, we show the plots for Hilbert-Schmidt distance [see Figs. 2(a) and 3(a)], quantum relative entropy [see Figs. 2(b) and 3(b)], and trace distance [see Figs. 2(c) and 3(c)], as functions of the dimensionless parameter $Jt$. The blue solid lines refer to ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}(\rho(t),{\texttt{R}_{1}})$, with $\rho(t)={e^{t\mathcal{L}}}[{\rho}({t_{0}})]$, while ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})$ is represented by red dashed lines, with ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]$.
The insets in Figs. 2 and 3 show the crossovers that clearly indicate the occurrence of the genuine Mpemba effect for each spin model. For the TFI model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.47$ for HSD [see Fig. 2(a)], (ii) $J{t_{\text{QME}}}\approx 6.16$ for QRE [see Fig. 2(b)], and (iii) $J{t_{\text{QME}}}\approx 1.57$ for TD [see Fig. 2(c)]. For the XXZ model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.23$ for HSD [see Fig. 3(a)], (ii) $J{t_{\text{QME}}}\approx 7.45$ for QRE [see Fig. 3(b), and (iii) $J{t_{\text{TD}}}\approx 1.85$ [see Fig. 3(c)]. The plots for the Hilbert-Schmidt distance and trace distance exhibit similar behavior when compared to each other, regardless of the spin model, except at the beginning of the nonunitary evolution. Indeed, the insets show that the HSD in Figs. 2(a) and 3(a) exhibit a more pronounced decay than the TD in Figs. 2(c) and 3(c), at earlier times of the dynamics, for both the TFI and XXZ models.
@@ VI Accelerated relaxation and quantum coherences
Let $\rho({t_{0}})$ an initial state of an open quantum system described by a Davies map. This state may or may not be incoherent in the energy eigenbasis. We recall that ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$, where ${\texttt{L}^{\prime}_{2}}=|{k_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian matrix that has a single nonzero element related to the pair $({k_{0}},{l_{0}})\in\{1,2,\ldots,d\}$, with ${k_{0}}\neq{l_{0}}$. Then, the overlap between $\texttt{L}_{2}$} and $\rho({t_{0}})$ becomes
@@
It is worth noting that Eq. (VI) allows for three possible scenarios. First, a vanishing overlap is obtained whenever $\rho({t_{0}})$ is an incoherent state in the energy eigenbasis. This means that the relaxation to equilibrium is expected to be accelerated. Second, for an initial state with nonzero coherences in this preferred basis, particularly with respect to the states $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, we obtain $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle\neq 0$. In this case, faster relaxation can be achieved by using the dressed state ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ as an input to the nonunitary evolution. Finally, there may exist a probe state $\rho({t_{0}})$ that exhibit nonzero coherences for all eigenstates of the Hamiltonian, except for the vector pair $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, i.e., $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle=0$. To see this, it is sufficient that the matrix ${U_{1}^{\dagger}}\rho({t_{0}}){U_{1}}$ has a vanishing off-diagonal element whose position coincides with that one of the single nonzero off-diagonal element of the eigenmatrix $\texttt{L}^{\prime}_{2}$.
@@
To illustrate these findings, we consider a system of two-qubits with Hamiltonian $H={U_{1}}\varepsilon{U_{1}^{\dagger}}$, with $\varepsilon=\text{diag}({\varepsilon_{1}},\ldots,{\varepsilon_{4}})$, where the energies are listed in descending order, ${\varepsilon_{j}}>{\varepsilon_{j+1}}$ for $j=\{1,2,3\}$. Here, ${U_{1}}=|{0,+}\rangle\langle{0,0}|+|{0,-}\rangle\langle{0,1}|+|{1,0}\rangle\langle{1,0}|+|{1,1}\rangle\langle{1,1}|$ is the unitary that brings the Hamiltonian of the system into its diagonal form, with $|{\pm}\rangle=(1/\sqrt{2})(|{0}\rangle\pm|{1}\rangle)$. Let ${\texttt{L}^{\prime}_{2}}=|{0,1}\rangle\langle{0,0}|$ be the eigenoperator related to the slowest decaying mode of the Liouvillian. On the one hand, for the initial state $\rho({t_{0}})=|{0,0}\rangle\langle{0,0}|$, we have that ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{0,+}\rangle\langle{0,+}|$, and one obtains the nonzero off-diagonal element $\langle{0,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{0,1}\rangle=1/2$. In this case, the dynamics of the quantum system can be exponentially accelerated by applying a suitable unitary matrix on the probe state such that it becomes diagonal respective to the eigenbasis of the Hamiltonian. In this case, to speed up the relaxation process, we follow our protocol and map $\rho({t_{0}})$ into the dressed state $\rho({t_{0}})$. On the other hand, when choosing the probe state $\rho({t_{0}})=|{1,+}\rangle\langle{1,+}|$, one obtains ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{1,+}\rangle\langle{1,+}|$, which implies a zero valued off-diagonal element $\langle{1,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{1,1}\rangle=0$. Hence, it follows that the dynamics of the quantum system is already exponentially accelerated, even thought $\rho({t_{0}})$ is not incoherent with respect to the energy eigenbasis.
@@ VII Conclusions
To characterize the relaxation process, we monitor the distinguishability between instantaneous states and the steady state of the open quantum system. To do so, we consider Hilbert-Schmidt distance, quantum relative entropy, and trace distance. We have shown that, for any initial state, there always exists a permutation matrix that maximizes the distance from the dressed state to equilibrium. Compared to the results of Ref. Moroder et al. (2024), our protocol is implementation-friendly since its numerical simulation requires low computational effort. We emphasize that our framework is capable of retrieving previous results discussed in the literature, making appropriate modifications and refining some assumptions. Our results contribute to elucidate the mechanism underlying the quantum Mpemba effect, and pave the way for insightful perspectives for the study of the dynamics of open quantum systems.
@@ A Exponentially accelerated relaxation for real slowest decaying mode
To be clear, the role of the permutation matrix $P_{\pi}$} is to rearrange the entries along the diagonal of the matrix $D$. In detail, given $D=\text{diag}({\lambda_{1}},{\lambda_{2}},\ldots,{\lambda_{d}})$, we have that ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, where $\pi(j)=k$ defines the permutation operation mapping a given number $j$ into $k$, with $j,k\in\{1,\ldots,d\}$.
@@
The right-hand side of Eq. (B13) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},{c_{2}},\ldots,{c_{d}}\}$, for all $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$. This means that there exists two permutation matrices $P_{\pi,\text{max}}$} and $P_{\pi,\text{min}}$} that satisfy the upper bounds $\text{Tr}({P_{\pi,\text{max}}}{D}{P_{\pi,\text{max}}^{\dagger}}\Sigma)\leq{c_{\text{max}}}$ and $\text{Tr}({P_{\pi,\text{min}}}{D}{P_{\pi,\text{min}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$, respectively. The inequality in Eq. (B13) is satisfied in both scenarios, but we realize that the latter case provides a tighter upper bound than the former. In particular, there must exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
@@
where $\pi(\bullet)$ maps index $j$ to $\pi(j)$, with $j,\pi(j)\in\{1,2,\ldots,d\}$. Therefore, to ensure that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},{\lambda_{\pi(2)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa. Indeed, it is know from the so-called rearrangement inequality that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}$ takes its minimum value if ${\alpha_{1}}\geq{\alpha_{2}}\geq\ldots\geq{\alpha_{d}}$ and ${\lambda_{\pi(1)}}\leq{\lambda_{\pi(2)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ Hardy et al. (1988). This proves the validity of the inequality in Eq. (B13), which also guarantees that Eq. (B1) is valid.

### v2

@@ V.2 Many-body quantum systems
Figure 4: (Color online) The quantum Mpemba effect in the dynamics of the XXZ model with open boundary conditions [see details in Sec. V.2], where $N=5$ spins, $\Delta=J/2$, $\gamma=1$, and $T=0.1$. To investigate the relaxation of the instantaneous states $\rho(t)$ and ${\rho^{\prime}}(t)$ towards the steady state $\texttt{R}_{1}$}, we consider the following figures of merit: (a) Hilbert-Schmidt distance, ${\mathcal{D}_{\text{HSD}}}(\rho(t),{\texttt{R}_{1}})$; (b) quantum relative entropy, ${\mathcal{D}_{\text{QRE}}}(\rho(t),{\texttt{R}_{1}})$; (c) trace distance, ${\mathcal{D}_{\text{TD}}}(\rho(t),{\texttt{R}_{1}})$. In each panel, we plot each of these distinguishability measures as a function of the dimensionless parameter $Jt$. The blue solid lines refer to the initial state $\rho({t_{0}})$ of the system, while the dressed state ${\rho^{\prime}}({t_{0}})$ is represented by red dashed lines.
Figures 3 and 4 present the numerical simulations for TFI and XXZ models, respectively. We set $N=5$, $h=J/2$, $\Delta=J/2$, $\gamma=1$, $T=0.1$, and ${k_{B}}=1$. In these figures, we show the plots for Hilbert-Schmidt distance [see Figs. 3(a) and 4(a)], quantum relative entropy [see Figs. 3(b) and 4(b)], and trace distance [see Figs. 3(c) and 4(c)], as functions of the dimensionless parameter $Jt$. The blue solid lines refer to ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}(\rho(t),{\texttt{R}_{1}})$, with $\rho(t)={e^{t\mathcal{L}}}[{\rho}({t_{0}})]$, while ${\mathcal{D}_{\text{HSD},\text{QRE},\text{TD}}}({\rho^{\prime}}(t),{\texttt{R}_{1}})$ is represented by red dashed lines, with ${\rho^{\prime}}(t)={e^{t\mathcal{L}}}[{\rho^{\prime}}({t_{0}})]$.
The insets in Figs. 3 and 4 show the crossovers that clearly indicate the occurrence of the genuine Mpemba effect for each spin model. For the TFI model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.47$ for HSD [see Fig. 3(a)], (ii) $J{t_{\text{QME}}}\approx 6.16$ for QRE [see Fig. 3(b)], and (iii) $J{t_{\text{QME}}}\approx 1.57$ for TD [see Fig. 3(c)]. For the XXZ model, we find the crossover to occur at (i) $J{t_{\text{QME}}}\approx 1.23$ for HSD [see Fig. 4(a)], (ii) $J{t_{\text{QME}}}\approx 7.45$ for QRE [see Fig. 4(b), and (iii) $J{t_{\text{TD}}}\approx 1.85$ [see Fig. 4(c)]. The plots for the Hilbert-Schmidt distance and trace distance exhibit similar behavior when compared to each other, regardless of the spin model, except at the beginning of the nonunitary evolution. Indeed, the insets show that the HSD in Figs. 3(a) and 4(a) exhibit a more pronounced decay than the TD in Figs. 3(c) and 4(c), at earlier times of the dynamics, for both the TFI and XXZ models.
@@ VI Accelerated relaxation and quantum coherences
Let $\rho({t_{0}})$ be an initial state of an open quantum system described by a Davies map. This state may or may not be incoherent in the energy eigenbasis. We recall that ${\texttt{L}_{2}}={U_{1}}{\texttt{L}^{\prime}_{2}}{U_{1}^{\dagger}}$, where ${\texttt{L}^{\prime}_{2}}=|{k_{0}}\rangle\langle{l_{0}}|$ is a non-Hermitian matrix that has a single nonzero element related to the pair $({k_{0}},{l_{0}})\in\{1,2,\ldots,d\}$, with ${k_{0}}\neq{l_{0}}$. Then, the overlap between $\texttt{L}_{2}$} and $\rho({t_{0}})$ becomes
@@
It is worth noting that Eq. (VI) allows for three possible scenarios. First, a vanishing overlap is obtained whenever $\rho({t_{0}})$ is an incoherent state in the energy eigenbasis. This means that the relaxation to equilibrium is expected to be accelerated. Second, for an initial state with nonzero coherences in this preferred basis, particularly with respect to the states $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, we obtain $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle\neq 0$. In this case, faster relaxation can be achieved by using the dressed state ${\rho^{\prime}}({t_{0}})=U\rho({t_{0}}){U^{\dagger}}$ as an input to the nonunitary evolution. Finally, there may exist a probe state $\rho({t_{0}})$ that exhibit nonzero coherences for all eigenstates of the Hamiltonian, except for the vector pair $|{l_{0}}\rangle$ and $|{k_{0}}\rangle$, i.e., $\text{Tr}({\texttt{L}_{2}}\rho({t_{0}}))=\langle{\psi_{l_{0}}}|\rho({t_{0}})|{\psi_{k_{0}}}\rangle=0$. To verify this, it is sufficient for matrix ${U_{1}^{\dagger}}\rho({t_{0}}){U_{1}}$ to have an off-diagonal entry equal to zero at the coordinate $({l_{0}},{k_{0}})$, while the only non-zero off-diagonal entry of matrix $\texttt{L}^{\prime}_{2}$} is located at $({k_{0}},{l_{0}})$.
@@
To illustrate these findings, we consider a system of two-qubits with Hamiltonian $H={U_{1}}\varepsilon{U_{1}^{\dagger}}$, with $\varepsilon=\text{diag}({\varepsilon_{1}},\ldots,{\varepsilon_{4}})$, where the energies are listed in descending order, ${\varepsilon_{j}}>{\varepsilon_{j+1}}$ for $j=\{1,2,3\}$. Here, ${U_{1}}=|{0,+}\rangle\langle{0,0}|+|{0,-}\rangle\langle{0,1}|+|{1,0}\rangle\langle{1,0}|+|{1,1}\rangle\langle{1,1}|$ is the unitary that brings the Hamiltonian of the system into its diagonal form, with $|{\pm}\rangle=(1/\sqrt{2})(|{0}\rangle\pm|{1}\rangle)$. Let ${\texttt{L}^{\prime}_{2}}=|{0,1}\rangle\langle{0,0}|$ be the eigenoperator related to the slowest decaying mode of the Liouvillian. On the one hand, for the initial state $\rho({t_{0}})=|{0,0}\rangle\langle{0,0}|$, we have that ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{0,+}\rangle\langle{0,+}|$, and one obtains the nonzero off-diagonal element $\langle{0,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{0,1}\rangle=1/2$. In this case, the dynamics of the quantum system can be accelerated exponentially by mapping ${\rho}({t_{0}})$ to the dressed state ${\rho^{\prime}}({t_{0}})$, according to our protocol. On the other hand, when choosing the probe state $\rho({t_{0}})=|{1,+}\rangle\langle{1,+}|$, one obtains ${U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}=|{1,+}\rangle\langle{1,+}|$, which implies a zero valued off-diagonal element $\langle{1,0}|{U_{1}}\rho({t_{0}}){U_{1}^{\dagger}}|{1,1}\rangle=0$. Hence, it follows that the dynamics of the quantum system is already exponentially accelerated, even thought $\rho({t_{0}})$ is not incoherent with respect to the energy eigenbasis.
@@ VII Conclusions
To characterize the relaxation process, we monitor the distinguishability between instantaneous states and the steady state of the open quantum system. To do so, we consider Hilbert-Schmidt distance, quantum relative entropy, and trace distance. We have shown that, for any initial state, there always exists a permutation matrix that maximizes the distance from the dressed state to equilibrium. In contrast to the results in Ref. Moroder et al. (2024), our protocol significantly reduces the computational cost of numerical simulations. We emphasize that our framework can reproduce the results of Ref. Carollo et al. (2021), with appropriate modifications and refinements to certain assumptions [see Appendix A]. Our results contribute to elucidate the mechanism underlying the quantum Mpemba effect, and pave the way for insightful perspectives for the study of the dynamics of open quantum systems.
@@ A Exponentially accelerated relaxation for real slowest decaying mode
To be clear, the role of the permutation matrix $P_{\pi}$} is to rearrange the entries along the diagonal of the matrix $D$. In detail, given $D=\text{diag}({\lambda_{1}},\ldots,{\lambda_{d}})$, we have that ${P_{\pi}}{D}{P_{\pi}^{\dagger}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$, where $\pi(j)=k$ defines the permutation operation mapping a given number $j$ into $k$, with $j,k\in\{1,\ldots,d\}$.
@@
The right-hand side of Eq. (B13) defines a convex sum ${\sum_{l}}\,{\epsilon_{l}}{c_{l}}$ of the nonnegative elements ${c_{l}}:=\text{Tr}({A_{l}}{D}{A_{l}^{\dagger}}\Sigma)$. We note that ${c_{\text{min}}}\leq{\sum_{l}}\,{\epsilon_{l}}{c_{l}}\leq{c_{\text{max}}}$, where ${c_{\text{min}}}=\text{min}\{{c_{1}},\ldots,{c_{d}}\}$ and ${c_{\text{max}}}=\text{max}\{{c_{1}},\ldots,{c_{d}}\}$, for all $0\leq{\epsilon_{l}}\leq 1$, with ${\sum_{l}}\,{\epsilon_{l}}=1$. This means that there exists two permutation matrices $P_{\pi,\text{max}}$} and $P_{\pi,\text{min}}$} that satisfy the upper bounds $\text{Tr}({P_{\pi,\text{max}}}{D}{P_{\pi,\text{max}}^{\dagger}}\Sigma)\leq{c_{\text{max}}}$ and $\text{Tr}({P_{\pi,\text{min}}}{D}{P_{\pi,\text{min}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$, respectively. The inequality in Eq. (B13) is satisfied in both scenarios, but we realize that the latter case provides a tighter upper bound than the former. In particular, there must exist an optimal permutation matrix $P_{\pi,\text{opt}}$} such that $\text{Tr}({P_{\pi,\text{opt}}}{D}{P_{\pi,\text{opt}}^{\dagger}}\Sigma)\leq{c_{\text{min}}}$. To see this point, we note that
@@
where $\pi(\bullet)$ maps index $j$ to $\pi(j)$, with $j,\pi(j)\in\{1,\ldots,d\}$. Therefore, to ensure that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}\leq{c_{\text{min}}}$, it suffices to choose a permutation matrix such that ${P_{\pi,\text{opt}}}{D}{P^{\dagger}_{\pi,\text{opt}}}=\text{diag}({\lambda_{\pi(1)}},\ldots,{\lambda_{\pi(d)}})$, with ${\lambda_{\pi(1)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ being listed in ascending order, provided that ${\alpha_{1}}\geq\ldots\geq{\alpha_{n}}$ are listed in descending order, or vice-versa. Indeed, it is know from the so-called rearrangement inequality that ${\sum_{j=1}^{d}}\,{\alpha_{j}}{\lambda_{\pi(j)}}$ takes its minimum value if ${\alpha_{1}}\geq\ldots\geq{\alpha_{d}}$ and ${\lambda_{\pi(1)}}\leq\ldots\leq{\lambda_{\pi(d)}}$ Hardy et al. (1988). This proves the validity of the inequality in Eq. (B13), which also guarantees that Eq. (B1) is valid.